OpenStudy (anonymous):

Select the equation of the line in slope-intercept form: 4x – 9y = –36 A. y = -4/9 x - 4 B. 9y = -4x - 36 C. 9y = 4x + 36 D. y=4/9 x + 4

OpenStudy (anonymous):

y=mx+c slope intercept form

OpenStudy (anonymous):

convert in it

OpenStudy (anonymous):

4x=-36+9y 9y=4x+36 y=4x/9+36/9 y=4x/9+4

OpenStudy (anonymous):

this